OpenStudy (falco276):

Help please? :( Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (falco276):

@Agent_Sniffles @Hope99 @e.mccormick @Compassionate @bahrom7893 @lakenclark @Loser66

OpenStudy (shamil98):

Where is my tag :( f(g(x)) = ( f o g ) f(x) = -4x + 7 and g(x) = 10x - 6 (f o g ) = -4 (10x-6) + 7 = -40x + 31

OpenStudy (shamil98):

you just plug in g(x) into f(x) :P

OpenStudy (falco276):

Sorry, forgot about you. *blushes.* Once again here's your reward. :)